OpenStudy (anonymous):

A box of 15 gadgets is known to contain 5 defective gadgets. If 4 gadgets are drawn at random, what is the probability of finding not more than 2 defective gadgets?

OpenStudy (anonymous):

@imqwerty

OpenStudy (anonymous):

Hey there. First, find p and q. Like I showed in your previous posted question. p is the probability of finding a defective and q is of not finding a defective one. p=5/15=1/3 q=1-p=1-1/3=2/3 np=15*1/3=5 nq=15*2/3=10 Both np and np are less than 5, hence a binomial expansion is to be used.\[P(X \le2)=P(X=0)+P(X=1)+P(X=2)\] Basing from the original formula\[P(X=x)=(nCx)*p^x*q ^{n-x}\] Try it out and tell me what you get.

OpenStudy (anonymous):

Any difficulties @Devanshi?

OpenStudy (anonymous):

Why did we multiply np and nq ?

OpenStudy (anonymous):

To know weither we use Binomial or Normal approximations to solve the problem.

OpenStudy (welshfella):

I knew it might be a binomial distribution bu I couldn't remember the details..

OpenStudy (anonymous):

If either or both np and nq are less than 5, then a Binomial Distribution is to be used.

OpenStudy (welshfella):

@Hoslos - don't you just check on the value of np? correct me if i'm wrong

OpenStudy (welshfella):

oh either / or OK

OpenStudy (anonymous):

@Devanshi ?